OpenStudy (tester97):

Use the change of base formula to evaluate log (base 7) 76.

OpenStudy (tester97):

\[\log_{7} 76\]

zepdrix (zepdrix):

Hmm I don't think 76 is any nice power... So we can use a calculator for this one, yes?

OpenStudy (tester97):

I get 64.22 when i do it with a calculator is that correct|?

zepdrix (zepdrix):

Here is the rule we want to use:\[\Large\bf\sf \color{orangered}{\log_a(b)\quad=\quad \frac{\log_{10}b}{\log_{10}a}}\]We'll change them to log base 10 so we can throw it into the calc.

zepdrix (zepdrix):

\[\Large\bf\sf \frac{\log76}{\log7}\quad\approx\quad ?\]

zepdrix (zepdrix):

Hmm are you sure you put it into the calculator correctly? You should get a very very different answer.

OpenStudy (tester97):

wait hold on let me try something else

zepdrix (zepdrix):

\[\Large\bf\sf \log(76)\div \log(7)\]Make sure you use brackets to keep things organized.

OpenStudy (tester97):

i get 2.22 is that correct?

zepdrix (zepdrix):

yayy good job \c:/

zepdrix (zepdrix):

Should round up to 2.23 I guess :P

OpenStudy (tester97):

Yes finally today i get something correct thank you.